# On heavy-flavour jets with Soft Drop

### 18 Dec 2023 v1

@@ 1 Introduction and motivation
Flavour physics has been studied for decades, both in the quark and lepton sectors. However, the recent development of Infra-Red and Collinear (IRC) safe flavour-jet algorithms Banfi:2006hf ; Caletti:2022hnc ; Czakon:2022wam ; Gauld:2022lem ; Caola:2023wpj opens up the possibility to set up a yet-unexplored flavour physics program that exploits jets and their substructure at the Large Hadron Collider (LHC).
@@ 2 Soft Drop
Soft Drop Larkoski:2014wba is a grooming algorithm that recursively removes soft-wide angle constitutes from a jet. The Soft Drop procedure starts by re-clustering a given jet (typically an anti-$k_{t}$ Cacciari:2008gp jet) with radius $R_{0}$ and transverse momentum $p_{t}$ with the Cambridge-Aachen (C/A) algorithm Wobisch:1998wt ; Dokshitzer:1997in . Soft Drop then parses the resulting angular-ordered branching history, grooming away the softer branch, until a branch that satisfies the condition
@@ 3.3 The $\theta_{g}$ distribution for a heavy-flavour jet
We now perform the resummation of the $\theta_{g}$ distribution for jets initiated by a heavy quark, considering both the case of a $b$-quark and a of $c$-quark. Both the Sudakov contribution $R$ and the clustering correction function $S$ in Eq. (4) need to be reconsidered. We start with the calculation of the resummed Sudakov exponent.
@@
$R_{i}(\theta_{g}^{2},m_{i}^{2},m_{b}^{2},m_{c}^{2})$ $=\int_{0}^{p_{t}^{2}R_{0}^{2}}\frac{\operatorname{d}\!k_{t}^{2}}{k_{t}^{2}+zm_{i}^{2}}\int_{0}^{1}\operatorname{d}\!z\,P_{gi}(z,k_{t}^{2})$ (14)
@@
where $\theta=\frac{k_{t}}{zp_{t}R_{0}}$. Note that the mass dependence enters in two ways. First, for a jet initiated by a quark flavour $i=b,c$, the quasi-collinear phase-space and the massive splitting function depend on $m_{i}$. Second, regardless of the jet flavour, the running of the strong coupling may cross the $b$ and the $c$ thresholds, thus inducting a logarithmic dependence on the quark masses. In order to proceed further, it is convenient to change the integration variable in Eq. (14) from $k_{t}^{2}$ to $\theta^{2}$:
@@
for $i=b,c$. The Sudakov exponent $R_{i}$ is given in Eq. (3.3) and the clustering contribution $S_{i}$ in Eq. (3.3). Note that our NLL distribution is no longer normalised, as a consequence of the fact that, in the massive case, the resummed cumulative behaves as a constant for $\theta_{g}\to 0$, while it vanishes in the massless case, see Eq. (4). Indeed, while in the massless case, Soft Drop with $\beta\geq 0$ behaves perturbatively as a groomer, i.e. within resumed perturbation theory, it always returns a jet with $\theta_{g}>0$, the quark mass provides an effective cutoff so that, there is a non-vanishing probability, given by
$S_{i}(0,\theta_{i}^{2},\xi_{b},\xi_{c})\,e^{-R_{i}(0,\theta_{i}^{2},\xi_{b},\xi_{c})}\sim e^{-\frac{\alpha_{\text{S}}C_{\text{F}}}{\pi\beta}\left(\log^{2}(z_{\text{cut}}\theta_{i}^{\beta})-\log^{2}(z_{\text{cut}})\right)},$ (31)
@@ 3.4 Numerical results and comparison to Monte Carlo simulations
We now provide numerical results for the $\theta_{g}$ distribution according to Eq. (30), and compare them to Monte Carlo (MC) simulations. To this purpose, we generate events using H ERWIG , version 7.2.2 Bellm:2019zci . To test the all-order behaviour of the observables we are interested in, we consider LO matrix elements for the hard process, dressed with the parton shower.  [^1] For detailed comparisons between analytic resummation and angular-ordered parton shower, we refer the reader to Hoang:2018zrp . Hadronisation effects are included using the H ERWIG cluster hadronisation model, when explicitly declared, and the CT14 Dulat:2015mca set of PDFs has been used throughout the paper.
@@
The $\theta_{g}$ distribution is shown in Fig. 2. Each plot includes parton level, i.e. with parton shower effect only, and hadron level, i.e. with hadronisation and the Underlying Event (UE) included. Together with the MC prediction, we show the corresponding NLL result of Eq. (30). The latter also exhibits a theoretical uncertainty given by the variation of the renormalisation scale, for which the central value is set at the hard scale $p_{t}R_{0}$, by a factor of two, as customary. Plots in different rows correspond to results for the identified leading jet in $Z+b$ ($B$), $Z+c$ ($D$) and $Z\,+$ light quark/hadron production, respectively. We always consider Soft Drop with $z_{\text{cut}}=0.1$, and $\beta=0$ and $1$, for left and right plots, respectively. For $b$ and $c$ jets, we also indicate, in green, the expected dead-cone region $\theta_{g}<\theta_{i}=\frac{m_{i}}{p_{t}R_{0}}$. All curves are normalised to have unit area. For $\beta=0$, i.e. when Soft Drop is more aggressive, we find good agreement between the NLL prediction and the MC. They are fairly different, instead, for $\beta=1$ at high $\theta_{g}$. However, in this region, we expect fixed-order corrections, not considered here, to be important.
@@ 4.1 Recap of the massless calculation
$p(z_{g}|\theta_{g})=\frac{P_{gq}(z_{g})\alpha_{\text{S}}(z_{g}^{2}\theta_{g}^{2}p_{t}^{2}R_{0}^{2})}{\int_{z_{\text{cut}}\theta_{g}^{\beta}}^{1}\operatorname{d}\!z\,P_{gq}(z)\alpha_{\text{S}}(z^{2}\,\theta_{g}^{2}p_{t}^{2}R_{0}^{2})}\Theta(z_{g}-z_{\text{cut}}\theta_{g}^{\beta}).$ (34)
@@
$R^{(\text{f.c.})}(\theta_{g}^{2})$ $=\frac{\alpha_{\text{S}}C_{\text{F}}}{\pi\beta}\Big{(}\log^{2}(z_{\text{cut}}\theta_{g}^{\beta})-\log^{2}(z_{\text{cut}})\Big{)}$
@@
$\frac{1}{\sigma_{0}}\frac{\operatorname{d}\!\sigma^{(\text{f.c.})}}{\operatorname{d}\!z_{g}}=P_{gq}(z_{g})\frac{\alpha_{\text{S}}}{2\pi}\int_{0}^{1}\frac{\operatorname{d}\!\theta_{g}^{2}}{\theta_{g}^{2}}e^{-R(\theta_{g}^{2})}\Theta(z_{g}-z_{\text{cut}}\theta_{g}^{\beta})$
@@
$\frac{1}{\sigma_{0}}\frac{\operatorname{d}\!\sigma^{(\text{f.c.})}}{\operatorname{d}\!z_{g}}=\frac{P_{gq}(z_{g})}{\int_{z_{c}}^{1}\operatorname{d}\!z\,P_{gq}(z)}\Theta(z_{g}>z_{\text{cut}})+{\cal O}\left(\alpha_{s}\right).$ (38)
It can be shown that the $\beta=0$ case does have a valid perturbative expansion in $\alpha_{\text{S}}$, despite being $\alpha_{\text{S}}$-independent at lowest order. Eq. (38) shows that the $\beta=0$ $z_{g}$-distribution is essentially driven by the QCD splitting function Larkoski:2015lea . This observation has initiated numerous theoretical Mehtar-Tani:2016aco ; KunnawalkamElayavalli:2017hxo ; Chang:2017gkt ; Milhano:2017nzm and experimental ATLAS:2019mgf ; ALargeIonColliderExperiment:2021mqf ; ALICE:2019ykw ; STAR:2021kjt ; CMS:2017qlm ; Kauder:2017cvz ; Kauder:2017mhg studies (see also Tripathee:2017ybi ; Larkoski:2017bvj ) that aim to use $z_{g}$ as probe QCD dynamics, both in $pp$ and heavy-ion collisions.
@@ 4.2 The $z_{g}$ distribution for a heavy-flavour jet
×$\times\Theta(z_{g}-z_{\text{cut}}\bar{\theta}^{\beta})\;\frac{\operatorname{d}\!\sigma_{i}}{\operatorname{d}\!\theta_{g}^{2}}\Bigg{|}_{\theta_{g}^{2}=\bar{\theta}^{2}-\theta_{i}^{2}},\quad i=b,c,$ (45)
@@
×$\times S_{i}(\bar{\theta}^{2},\theta_{i}^{2},\xi_{b},\xi_{c})\,e^{-R_{i}(\bar{\theta}^{2},\theta_{i}^{2},\xi_{b},\xi_{c})}\Theta(z_{g}-z_{\text{cut}}\bar{\theta}^{\beta}).$ (46)
@@ 4.3 Numerical results and comparison to Monte Carlo simulations
Figure 4: The groomed momentum fraction ($z_{g}$) distribution for $b$-jets (top), $c$-jets (middle) and light-quark jets (bottom). In all plots, we show our resumed result, as well as the one obtained with the Monte Carlo event generator H
@@
, both at parton and hadron level. Jets are selected with the anti-$k_{t}$ algorithm with $R_{0}=0.4$, and groomed with Soft Drop with $z_{c}=0.1$ and $\beta=0$. The plots on the left are for $p_{t}\geq 50$ GeV, while the ones on the left are for $p_{t}\geq 150$ GeV. The uncertainty bands for the analytic predictions are obtained by varying the resummation scale by a factor of two above and below the hard scale $p_{t}R_{0}$, i.e. $\mu_{R}\in\left[\frac{p_{t}R_{0}}{2},2p_{t}R_{0}\right]$.
@@
In Fig. 4, we compare resummed results for two different values of the jet transverse momentum cut, namely $p_{t}\geq 50$ GeV, on the left, and $p_{t}\geq 150$ GeV, on the right. The plots at the top are for $b$-jets, the ones in the middle for $c$-jets, and the ones at the bottom for light-quark jets. As already pointed out, the results all look very similar, because the shape of the distribution is mostly driven by the singular part of the splitting function for the emission of a gluon off a (massive) quark.
Overall, we find good agreement between our calculation and the Monte Carlo results, although our results are strangely closer to the full simulation than to the parton-level one. We note that our predictions undershoot the Monte Carlo at large $z_{g}$. We have traced this back to the fact that our calculation does not take into account symmetrised splitting functions, as discussed in the footnote of Sec. 4.1.
@@ 4.4 Comparison to fragmentation functions
In what follows, we would like to study the correlation, if any, between the variable $\zeta$, which measures the departure from Born kinematics of the identified heavy flavour within a jet, to $z_{g}$ (with $\beta=0$), which measures probes the kinematics of the splitting that passes Soft Drop, in a heavy-flavour tagged jet.
@@ 5 Conclusions and Outlook
We have also considered the momentum fraction $z_{g}$ of the first emission that passes Soft Drop. Our calculation generalises to the massive case the conditional-probability approach of Larkoski:2015lea and it allows for a resummation of mass effects. We have compared our numerical results (for $\beta=0$) to Monte Carlo simulations. As for the massless case, the result is driven by the QCD splitting function and it is largely insensitive to non-perturbative effects.
Finally, we have compared the $z_{g}$ distribution (for $\beta=0$) to the fragmentation variable $x$, which measures the ratio of the heavy quark (hadron) transverse momentum to the jet $p_{t}$. We performed the analytic calculation of both distributions at $\mathcal{O}(\alpha_{\text{S}})$, showing that they lead to the same results for $z_{g}=1-x>z_{c}$. We have investigated the all-order behaviour of these observables using Monte Carlo simulations and discovered that the parton-shower significantly dilutes this correlation. However, we have found that, in contrast to $z_{g}$, the $x$ receives sizeable non-perturbative corrections. In the future, it would be interesting to repeat these studies with the modified version of the declustering procedure Cunqueiro:2018jbh , whereby one follows the flavoured branch.

### 29 Feb 2024 v2

@@ 1 Introduction and motivation
Flavour physics has been studied for decades, both in the quark and lepton sectors. However, the recent development of Infra-Red and Collinear (IRC) safe flavour-jet algorithms Banfi:2006hf ; Caletti:2022hnc ; Czakon:2022wam ; Gauld:2022lem ; Caola:2023wpj opens up the possibility of setting up a yet-unexplored flavour physics program that exploits jets and their substructure at the Large Hadron Collider (LHC).
@@ 2 Soft Drop
Soft Drop Larkoski:2014wba is a grooming algorithm that recursively removes soft-wide angle constituents from a jet. The Soft Drop procedure starts by re-clustering a given jet (typically an anti-$k_{t}$ Cacciari:2008gp jet) with radius $R_{0}$ and transverse momentum $p_{t}$ with the Cambridge-Aachen (C/A) algorithm Wobisch:1998wt ; Dokshitzer:1997in . Soft Drop then parses the resulting angular-ordered branching history, grooming away the softer branch, until a branch that satisfies the condition
@@ 3.3 The $\theta_{g}$ distribution for a heavy-flavour jet
We now perform the resummation of the $\theta_{g}$ distribution for jets initiated by a heavy quark, considering both the case of a $b$-quark and a $c$-quark. Both the Sudakov contribution $R$ and the clustering correction function $S$ in Eq. (4) need to be reconsidered. We start with the calculation of the resummed Sudakov exponent.
@@
$R_{i}(\theta_{g}^{2},m_{i}^{2},m_{b}^{2},m_{c}^{2})$ $=\int_{0}^{p_{t}^{2}R_{0}^{2}}\frac{\operatorname{d}\!k_{t}^{2}}{k_{t}^{2}+z^{2}m_{i}^{2}}\int_{0}^{1}\operatorname{d}\!z\,P_{gi}(z,k_{t}^{2})$ (14)
@@
where $\theta=\frac{k_{t}}{zp_{t}R_{0}}$. Note that the mass dependence enters in two ways. First, for a jet initiated by a quark flavour $i=b,c$, the quasi-collinear phase-space and the massive splitting function depend on $m_{i}$. Second, regardless of the jet flavour, the running of the strong coupling may cross the $b$ and the $c$ thresholds, thus inducing a logarithmic dependence on the quark masses. In order to proceed further, it is convenient to change the integration variable in Eq. (14) from $k_{t}^{2}$ to $\theta^{2}$:
@@
for $i=b,c$. The Sudakov exponent $R_{i}$ is given in Eq. (3.3) and the clustering contribution $S_{i}$ in Eq. (3.3). Note that our NLL distribution is no longer normalised, as a consequence of the fact that, in the massive case, the resummed cumulative behaves as a constant for $\theta_{g}\to 0$, while it vanishes in the massless case, see Eq. (4). Indeed, while in the massless case, Soft Drop with $\beta\geq 0$ behaves perturbatively as a groomer, i.e. within resummed perturbation theory, it always returns a jet with $\theta_{g}>0$, the quark mass provides an effective cutoff so that, there is a non-vanishing probability, given by
$S_{i}(0,\theta_{i}^{2},\xi_{b},\xi_{c})\,e^{-R_{i}(0,\theta_{i}^{2},\xi_{b},\xi_{c})}\sim e^{-\frac{\alpha_{\text{S}}C_{\text{F}}}{\pi\beta}\left(\log^{2}(z_{c}\theta_{i}^{\beta})-\log^{2}(z_{c})\right)},$ (31)
@@ 3.4 Numerical results and comparison to Monte Carlo simulations
We now provide numerical results for the $\theta_{g}$ distribution according to Eq. (30), and compare them to Monte Carlo (MC) simulations. To this purpose, we generate events using H ERWIG , version 7.2.2 Bellm:2019zci . To test the all-order behaviour of the observables we are interested in, we consider LO matrix elements for the hard process, dressed with the parton shower.  [^1] In this paper we concentrate on an angular-ordered parton shower, see Hoang:2018zrp for detailed comparisons with analytic resummation. We have also tested our analytic predictions against the H ERWIG dipole shower Platzer:2011bc and the P YTHIA 8.3 Bierlich:2022pfr one and we did not find significant differences. It would be interesting, in the future, to also compare to state-of-the-art simulations that also include matching to fixed order. Hadronisation effects are included using the H ERWIG cluster hadronisation model, when explicitly declared, and the CT14 Dulat:2015mca set of PDFs has been used throughout the paper.
@@
The $\theta_{g}$ distribution is shown in Fig. 2. Each plot includes parton level, i.e. with parton shower effect only, and hadron level, i.e. with hadronisation and the Underlying Event (UE) included. Together with the MC prediction, we show the corresponding NLL result of Eq. (30). The latter also exhibits a theoretical uncertainty given by the variation of the renormalisation scale, for which the central value is set at the hard scale $p_{t}R_{0}$, by a factor of two, as customary. Plots in different rows correspond to results for the identified leading jet in $Z+b$ ($B$), $Z+c$ ($D$) and $Z\,+$ light quark/hadron production, respectively. We always consider Soft Drop with $z_{c}=0.1$, and $\beta=0$ and $1$, for left and right plots, respectively. For $b$ and $c$ jets, we also indicate, in green, the expected dead-cone region $\theta_{g}<\theta_{i}=\frac{m_{i}}{p_{t}R_{0}}$. All curves are normalised to have unit area. For $\beta=0$, i.e. when Soft Drop is more aggressive, we find good agreement between the NLL prediction and the MC. They are fairly different, instead, for $\beta=1$ at high $\theta_{g}$. However, in this region, we expect fixed-order corrections, not considered here, to be important.
@@ 4.1 Recap of the massless calculation
$p(z_{g}|\theta_{g})=\frac{P_{gq}(z_{g})\alpha_{\text{S}}(z_{g}^{2}\theta_{g}^{2}p_{t}^{2}R_{0}^{2})}{\int_{z_{c}\theta_{g}^{\beta}}^{1}\operatorname{d}\!z\,P_{gq}(z)\alpha_{\text{S}}(z^{2}\,\theta_{g}^{2}p_{t}^{2}R_{0}^{2})}\Theta(z_{g}-z_{c}\theta_{g}^{\beta}).$ (34)
@@
$R^{(\text{f.c.})}(\theta_{g}^{2})$ $=\frac{\alpha_{\text{S}}C_{\text{F}}}{\pi\beta}\Big{(}\log^{2}(z_{c}\theta_{g}^{\beta})-\log^{2}(z_{c})\Big{)}$
@@
$\frac{1}{\sigma_{0}}\frac{\operatorname{d}\!\sigma^{(\text{f.c.})}}{\operatorname{d}\!z_{g}}=P_{gq}(z_{g})\frac{\alpha_{\text{S}}}{2\pi}\int_{0}^{1}\frac{\operatorname{d}\!\theta_{g}^{2}}{\theta_{g}^{2}}e^{-R(\theta_{g}^{2})}\Theta(z_{g}-z_{c}\theta_{g}^{\beta})$
@@
$\frac{1}{\sigma_{0}}\frac{\operatorname{d}\!\sigma^{(\text{f.c.})}}{\operatorname{d}\!z_{g}}=\frac{P_{gq}(z_{g})}{\int_{z_{c}}^{1}\operatorname{d}\!z\,P_{gq}(z)}\Theta(z_{g}>z_{c})+{\cal O}\left(\alpha_{s}\right).$ (38)
It can be shown that the $\beta=0$ case does have a valid perturbative expansion in $\alpha_{\text{S}}$, despite being $\alpha_{\text{S}}$-independent at lowest order. Eq. (38) shows that the $\beta=0$ $z_{g}$-distribution is essentially driven by the QCD splitting function Larkoski:2015lea . This observation has initiated numerous theoretical Mehtar-Tani:2016aco ; KunnawalkamElayavalli:2017hxo ; Chang:2017gkt ; Milhano:2017nzm and experimental ATLAS:2019mgf ; ALargeIonColliderExperiment:2021mqf ; ALICE:2019ykw ; STAR:2021kjt ; CMS:2017qlm ; Kauder:2017cvz ; Kauder:2017mhg studies (see also Tripathee:2017ybi ; Larkoski:2017bvj ) that aim to use $z_{g}$ as a probe of QCD dynamics, both in $pp$ and heavy-ion collisions.
@@ 4.2 The $z_{g}$ distribution for a heavy-flavour jet
×$\times\Theta(z_{g}-z_{c}\bar{\theta}^{\beta})\;\frac{\operatorname{d}\!\sigma_{i}}{\operatorname{d}\!\theta_{g}^{2}}\Bigg{|}_{\theta_{g}^{2}=\bar{\theta}^{2}-\theta_{i}^{2}},\quad i=b,c,$ (45)
@@
×$\times S_{i}(\bar{\theta}^{2},\theta_{i}^{2},\xi_{b},\xi_{c})\,e^{-R_{i}(\bar{\theta}^{2},\theta_{i}^{2},\xi_{b},\xi_{c})}\Theta(z_{g}-z_{c}\bar{\theta}^{\beta}).$ (46)
@@ 4.3 Numerical results and comparison to Monte Carlo simulations
Figure 4: The groomed momentum fraction ($z_{g}$) distribution for $b$-jets (top), $c$-jets (middle) and light-quark jets (bottom). In all plots, we show our resummed result, as well as the one obtained with the Monte Carlo event generator H
@@
, both at parton and hadron level. Jets are selected with the anti-$k_{t}$ algorithm with $R_{0}=0.4$, and groomed with Soft Drop with $z_{c}=0.1$ and $\beta=0$. The plots on the left are for $p_{t}\geq 50$ GeV, while the ones on the right are for $p_{t}\geq 150$ GeV. The uncertainty bands for the analytic predictions are obtained by varying the resummation scale by a factor of two above and below the hard scale $p_{t}R_{0}$, i.e. $\mu_{R}\in\left[\frac{p_{t}R_{0}}{2},2p_{t}R_{0}\right]$.
@@
In Fig. 4, we compare resummed results for two different values of the jet transverse momentum cut, namely $p_{t}\geq 50$ GeV, on the left, and $p_{t}\geq 150$ GeV, on the right. The plots at the top are for $b$-jets, the ones in the middle for $c$-jets, and the ones at the bottom for light-quark jets. All curves are normalised to have unit area. As already pointed out, the results all look very similar, because the shape of the distribution is mostly driven by the singular part of the splitting function for the emission of a gluon off a (massive) quark. However, we also note that differences between $b$,$c$, and light flavours are larger in the Monte Carlo results than in the analytic ones. This could be due to quark masses influencing the kinematics, thus causing mass-dependent recoil effects. These are, to a certain extent, accounted for in the parton shower, but neglected in the analytic resummation.
Overall, we find good agreement between our calculation and the Monte Carlo results, although our results are strangely closer to the full simulation than to the parton-level one. We note that our predictions undershoot the Monte Carlo at large $z_{g}$. We have traced this back to the fact that our calculation does not take into account symmetrised splitting functions, as discussed in the footnote of Sec. 4.1. Indeed by rescaling the $z_{g}$ distribution by the symmetrised splitting function, the tail of the analytic calculation moves closer to the Monte Carlo results.
@@ 4.4 Comparison to fragmentation functions
In what follows, we would like to study the correlation, if any, between the variable $\zeta$, which measures the departure from Born kinematics of the identified heavy flavour within a jet, to $z_{g}$ (with $\beta=0$), which probes the kinematics of the splitting that passes Soft Drop, in a heavy-flavour tagged jet.
@@ 5 Conclusions and Outlook
We have also considered the momentum fraction $z_{g}$ of the first emission that passes Soft Drop. Our calculation generalises to the massive case the conditional-probability approach of Larkoski:2015lea and it allows for a resummation of mass effects. While our results hold for any $\beta\geq 0$, we have focussed our numerical investigations on the $\beta=0$ case. This choice is motivated by a few reasons. First, this is the value for which, at least to lowest order, one has a clear factorisation of the conditional probability expression, leading to a distribution proportional to the splitting function. For $\beta>0$, the dependence on the splitting function is smeared out. Second, it is for the $\beta=0$ case that one expects to find more similarities to the fragmentation variable $x$, and third, from the study of the $\theta_{g}$ distribution, we have found that the $\beta=0$ is under better theoretical control and more sensitive to dead cone effects. We have compared our numerical results to Monte Carlo simulations. As for the massless case, the result is driven by the QCD splitting function and it is largely insensitive to non-perturbative effects.
Finally, we have compared the $z_{g}$ distribution to the fragmentation variable $x$, which measures the ratio of the heavy quark (hadron) transverse momentum to the jet $p_{t}$. We performed the analytic calculation of both distributions at $\mathcal{O}(\alpha_{\text{S}})$, showing that they lead to the same results for $z_{g}=1-x>z_{c}$. We have investigated the all-order behaviour of these observables using Monte Carlo simulations and discovered that the parton-shower significantly dilutes this correlation. However, we have found that, in contrast to $z_{g}$, the $x$ receives sizeable non-perturbative corrections. In the future, it would be interesting to repeat these studies with the modified version of the declustering procedure Cunqueiro:2018jbh , whereby one follows the flavoured branch.